OpenStudy (anonymous):

Find all real solutions of the equation. Show work: \[x^5 + 8x^2 =0\]

OpenStudy (anonymous):

I can get to: \[x^2(x^3+8)\] and I get stuck....

OpenStudy (anonymous):

8=\[2^3\]\[x^2(x^3+2^3)\]\[x^3+2^3=0\]\[x^3=-2^3\]\[x=-2\]\[x^2=0 ------x=0\]

OpenStudy (anonymous):

thank you. couldn't figure out how to factor that...

OpenStudy (anonymous):

welcome anytime